A news company reported that 6% of adult Americans have a food allergy. Consider selecting 15 adult Americans at random. Define the random variable x as

x = number of people in the sample of 15 that have a food allergy.

Find the following probabilities. (Round your answers to three decimal places.)


(a)

P(x < 3)


(b)

P(x ≤ 3)


(c)

P(x ≥ 4)


(d)

P(1 ≤ x ≤ 3)



1
Expert's answer
2021-12-06T11:40:30-0500

"x\sim Bin(n, p)"

Given "n=15, p=0.06, q=1-p=1-0.06=0.94."

(a)


"P(x<3)=P(x=0)+P(x=1)"

"+P(x=2)=\dbinom{15}{0}(0.06)^0(0.94)^{15-0}"

"+\dbinom{15}{1}(0.06)^1(0.94)^{15-1}+\dbinom{15}{2}(0.06)^2(0.94)^{15-2}"

"\approx0.3953+0.3785+0.1691\approx0.943"

(b)


"P(x\leq 3)=P(x<3)+P(x=3)"

"\approx0.9429+\dbinom{15}{3}(0.06)^3(0.94)^{15-3}"

"\approx0.9429+0.0468\approx0.990"

(c)


"P(x\geq 4)=1-P(x\leq 3)"

"\approx1-0.9897\approx0.010"

(d)


"P(1\leq x\leq 3)=P(x=1)+P(x=2)"

"+P(x=3)=\dbinom{15}{1}(0.06)^1(0.94)^{15-1}"

"+\dbinom{15}{2}(0.06)^2(0.94)^{15-2}+\dbinom{15}{3}(0.06)^2(0.94)^{15-3}"

"\approx0.3785+0.1691+0.0468\approx0.594"